OpenStudy (wondermath):

Solve the trig equation sin2x=1/3

OpenStudy (akshay_budhkar):

sin2x=1/3 2x=sin^-1(1/3) x=[sin^-1(1/3]/2

OpenStudy (anonymous):

are you sure ?

OpenStudy (akshay_budhkar):

yes

OpenStudy (wondermath):

yeah i got that, but there is more than one answer

OpenStudy (anonymous):

sin2x=2sinxcos which is a trig identity

OpenStudy (akshay_budhkar):

yea the other answer is sin2x=1/3 2x=(pie)-sin^-1(1/3)

OpenStudy (akshay_budhkar):

um as sinx takes the value twice in 0 to pie..

OpenStudy (akshay_budhkar):

you get it?

OpenStudy (akshay_budhkar):

@salem yes correct.. so what is x using that formula?

OpenStudy (wondermath):

actually in the back it has 4 answers

OpenStudy (wondermath):

i got the firs two

OpenStudy (akshay_budhkar):

4 answers lol

OpenStudy (wondermath):

lol yea

OpenStudy (akshay_budhkar):

actually it has infinite solutions

OpenStudy (wondermath):

LOL

OpenStudy (akshay_budhkar):

sinx is a continuous curve

OpenStudy (wondermath):

no sorry forgot to say its from 0 to 2pi

OpenStudy (anonymous):

thanks for correction Akshay

OpenStudy (akshay_budhkar):

umm lol lemme think i have not done trigo for a while.. u gotta use the formulae like sin 2(pie) + x=sinx and stuff

OpenStudy (anonymous):

you mean getting the function back to its the original for or use trig identity for angles

OpenStudy (akshay_budhkar):

sorry wondermath my brain needs to dig it out now lol .. i will ask some1 to help u :P do u get the concept?

OpenStudy (wondermath):

lol okay thanks